Evaluate the integral of 3 dx dy. Use inner limits=y^2 to 9, outer limits 0 to 3.



Determine the integral of xy dx dy. Using limits of x=0 to (y+1) and y=0 to 1.



Find the integral of (2-x-y) dx dy. Using limits of x=0 to (1-y)^0.5, and y=0 to 1.

1
Expert's answer
2021-12-20T17:08:01-0500

a)


"\displaystyle\int_{0}^3\displaystyle\int_{y^2}^{9}3dxdy=3\displaystyle\int_{0}^3[x]\begin{matrix}
 9 \\
 y^2
\end{matrix}dy"

"=3\displaystyle\int_{0}^3(9-y^2)dy=[27y-y^3]\begin{matrix}
 3 \\
 0
\end{matrix}"

"=81-27-0=54"

b)


"\displaystyle\int_{0}^1\displaystyle\int_{0}^{y+1}xydxdy=\displaystyle\int_{0}^{1}y[\dfrac{x^2}{2}]\begin{matrix}
 y+1 \\
 0
\end{matrix}dy"

"=\dfrac{1}{2}\displaystyle\int_{0}^1(y^3+2y^2+y)dy=\dfrac{1}{2}[\dfrac{y^4}{4}+\dfrac{2y^3}{3}+\dfrac{y^2}{2}]\begin{matrix}
 1 \\
 0
\end{matrix}"

"=\dfrac{1}{2}(\dfrac{1}{4}+\dfrac{2}{3}+\dfrac{1}{2})=\dfrac{17}{24}"

c)


"\displaystyle\int_{0}^1\displaystyle\int_{0}^{\sqrt{1-y}}(2-x-y)dxdy"

"=\displaystyle\int_{0}^{1}[2x-\dfrac{x^2}{2}-xy]\begin{matrix}
 \sqrt{1-y} \\
 0
\end{matrix}dy"

"=\displaystyle\int_{0}^{1}(2\sqrt{1-y}-\dfrac{1-y}{2}-y\sqrt{1-y})dy"

"=[-\dfrac{4(1-y)^{3/2}}{3}-\dfrac{y}{2}+\dfrac{y^2}{4}-\dfrac{2(1-y)^{5/2}}{5}+\dfrac{2(1-y)^{3/2}}{3}]\begin{matrix}
 1 \\
 0
\end{matrix}"

"=-\dfrac{1}{2}+\dfrac{1}{4}+\dfrac{2}{3}+\dfrac{2}{5}=\dfrac{49}{60}"